OpenStudy (anonymous):

Which expression is equivalent to (5^3)^−2?

OpenStudy (anonymous):

@bibby @BassCatcher15 @brala00 @nick67012

OpenStudy (jdoe0001):

\(\bf a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\qquad thus \\ \quad \\ (5^3)^{-{\color{red}{ 2}}}\implies ?\)

OpenStudy (anonymous):

http://prntscr.com/4jfmuc answer choices

OpenStudy (anonymous):

ok well first you must remember if you have a negative exponent you must flip the sign.\[(5^3)^-2 = (\frac{ 1 }{ 5^3 })^2\]\[\frac{ 1 }{ 5^6 }\]

OpenStudy (anonymous):

what does 5^6 equal too?

OpenStudy (anonymous):

15? @FriedRice

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Or is it 30?

OpenStudy (anonymous):

what is 5*5*5*5*5*5 =?

OpenStudy (anonymous):

15 625

OpenStudy (anonymous):

@friedrice

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the answer is 1/ 15625